# Guiding LLMs to Generate High-Fidelity and High-Quality Counterfactual Explanations for Text Classification

### v1

@@ Abstract
The need for interpretability in deep learning has driven interest in counterfactual explanations, which identify minimal changes to an instance that change a model’s prediction. Current counterfactual (CF) generation methods require task-specific fine-tuning and produce low-quality text. Large Language Models (LLMs), though effective for high-quality text generation, struggle with label-flipping counterfactuals (i.e., counterfactuals that change the prediction) without fine-tuning. We introduce two simple classifier-guided approaches to support counterfactual generation by LLMs, eliminating the need for fine-tuning while preserving the strengths of LLMs. Despite their simplicity, our methods outperform state-of-the-art counterfactual generation methods and are effective across different LLMs, highlighting the benefits of guiding counterfactual generation by LLMs with classifier information. We further show that data augmentation by our generated CFs can improve a classifier’s robustness. Our analysis reveals a critical issue in counterfactual generation by LLMs: LLMs rely on parametric knowledge rather than faithfully following the classifier. [^1] Source code is available at https://anonymous.4open.science/r/llm-cls-EC84/
@@ 1 Introduction
Counterfactual explanations identify minimal changes to the input that result in a changed prediction by a classifier. As such, they are a core tool for eXplainable AI (XAI), providing valuable insights into a classifier’s decision-making process [21, 31], thereby fostering interpretability. Counterfactual explanations in an XAI context differ from real-world counterfactuals: the latter describe hypothetical scenarios that contradict known facts [16], while the former aim to uncover a classifier’s decision-making process. For example, given a negatively classified movie review, identifying minimal changes (adding, replacing or deleting words) that would turn the classifier’s prediction into positive, while maintaining meaningful content (cf. Fig. 1) can provide insight into the classifier’s reasoning and decision boundary. This insight can expose undesired model behavior, such as biases against minorities [20], reliance on spurious correlations [11], or vulnerability to adversarial manipulations [12].
Counterfactual generation methods that access the classifier produce counterfactuals (CFs) with a high label-flip rate (i.e., they succeed in actually changing the prediction) and minimal changes [27, 29, 26]. However, they require task-specific fine-tuning and may generate low-quality text [22]. In contrast, Large Language Models (LLMs) are used for end-to-end generation without expensive fine-tuning and produce high-quality text [24, 15]. However, they struggle to flip the labels with minimal changes [22, 23], i.e., do not generate high-fidelity [^2] The terms fidelity and faithful are sometimes used synonymously to describe how well an explanation aligns with the true reasoning of a classifier. In this paper, we use fidelity to refer to “true counterfactuals”, i.e., counterfactuals that actually change a classifier’s prediction and faithful to describe whether a counterfactual truly explains the classifier’s actual reasoning. CFs, because they do not have information about the classifier and its decision boundary.
@@
In this paper, we aim to bridge the gap between fully supervised, costly CF generation methods and unsupervised, end-to-end text generation using LLMs. We present two simple yet effective approaches for incorporating classifier information to guide LLMs in CF generation (cf. Fig. 1), enabling high-quality text generation without the need for fine-tuning and offering broad applicability. In an ante-hoc approach, we first identify important words using an XAI method (saliency maps [28] and SHAP [18]). We then use the most important words and examples of factual-counterfactual pairs to prompt the LLM to generate counterfactual texts. Our second approach prompts the LLM to generate a set of candidate counterfactuals, evaluates their fidelity to the classifier, and selects the best candidate post-hoc based on minimality of changes and fidelity.
@@ 2 Related work
Counterfactual text generation methods fall into three categories [22]: masking and filling approaches, sampling based on conditional distributions, and LLM-based generation.
@@
These methods involve (1) masking key words and (2) filling the masked words using a pretrained language model. MICE [27] and AutoCAD [30] use classifier gradients to identify important words, while DoCoGen [4] masks n-grams with scores exceeding a threshold. CREST [29] trains SPECTRA [13] to detect masking candidates. In step (2), these methods fine-tune T5 [25] to fill the blanks. Polyjuice [31] allows user-defined masking and fine-tunes a RoBERTa-based model using control codes to fill the blanks.
@@
CF-GAN [26] and CLOSS [9] model a conditional distribution to generate counterfactuals based on the target label. CF-GAN adapts StarGAN [7] to produce realistic counterfactuals. CLOSS generates text counterfactuals by optimizing token embeddings, ranking impactful changes with Shapley values, and using beam search to select minimal, grammatically plausible edits that flip a model’s prediction.
@@
CORE [8] trains a counterfactual retriever using contrastive learning, where positive pairs consist of human-authored counterfactuals and negative pairs are paraphrases. The retriever selects relevant counterfactual excerpts from an unlabeled corpus, which are then incorporated into prompts to guide GPT-3 [3] in generating diverse counterfactual edits. CORE relies on counterfactual data and does not use classifier information. In contrast, DISCO [6] generates a large set of perturbations by masking spans identified by a neural syntactic parser and filling them using GPT-3. A filtering model then selects the most promising counterfactuals. However, DISCO is computationally expensive and therefore primarily suited for short-text tasks such as natural language inference (NLI). Additionally, DISCO focuses on generating CFs for data augmentation rather than explanations. FLARE [1] generates CFs by prompting LLMs in three steps: extracting latent features, identifying relevant words linked to those features, and minimally modifying these words to produce CFs.
In summary, MF and CD methods require task-specific fine-tuning, while LLM-based generation lacks classifier integration and produces poor fidelity counterfactuals [22]. Our work not only addresses the limitations of both, fine-tuned and LLM-based methods, but also combines their strengths.
@@ 3.1 Classifier-Guided Generation (CGG)
In CGG, the generation process is guided by information from the classifier. Specifically, we use an XAI method $X(f(x))$ to extract words from the input $x$ that are important to the classifier’s decision $f(x)$. These important words are then incorporated into the predefined prompt $p$ to form $p(x,X(f(x)))$ (see Fig. 2(a)). A complete example of the prompt is provided in Appendix 0.A. This prompt is input to the LLM $M$, which generates a counterfactual instance $x^{\prime}$, such that $M(p(x,X(f(x))))=x^{\prime}$.
@@ 3.2 Classifier-Guided Validation (CGV)
CGV uses the classifier after the generation process. Specifically, we use the LLM $M$ with prompt $p$ to generate $n$ candidate samples $\{\hat{x}_{i}\}_{i=1}^{n}$. The classifier $f$ is then used to validate the labels, assigning a label $\hat{y}_{i}$ to each sample, and we select the sample with the flipped target label $y^{\prime}$ that also minimizes the distance $d(\hat{x}_{i},x)$ to the original input $x$ (see Fig. 2(b)). Formally, we have
@@ 4 Experimental Setup
We evaluate counterfactuals on the CEVAL benchmark [22], which includes two datasets: IMDB [19], a sentiment analysis dataset with binary labels (positive and negative), and SNLI [2], a natural language inference dataset with three labels (entailment, contradiction, and neutral). The benchmark assesses performance by counterfactual and text quality metrics. A benchmark run required approximately two hours per dataset on an A100 80GB GPU.
@@ 4.1 Models
Following the benchmark [22], we use BERT-based classifiers $f$ that achieve 89% accuracy on IMDB and 90% on SNLI. We compare against MICE [27], the top method for counterfactual metrics according to the benchmark, CREST [29], a masking and filling method that uses SPECTRA [13] for masking and T5 [25] for filling in the blanks, and FLARE [1], an LLM-based method that generates counterfactuals through three steps by prompting LLMs. The baseline is vanilla 1-shot prompting without classifier guidance.
For the counterfactual generation LLM $M$, we compare Llama-3.1-8B-Instruct, Llama-2-7b-chat, and GPT-4o-mini [^3] Version: GPT-4o-mini-2024-07-18.. Following [22], we use a temperature of 1.0 for all LLMs in the counterfactual generation task. For CGG, we adopt saliency maps [28] as the XAI method $X$, selecting the top 25% most important words with the highest gradients relative to the classifier’s prediction, consistent with the benchmark (additional results with SHAP [18] are similar, cf. Appendix 0.B). For CGV, we use Levenshtein Distance as distance measure $d$ to select $x^{\prime}$ and set the number of samples $n=5$ for IMDB and $n=10$ for SNLI, as larger $n$ values showed minimal differences in preliminary experiments. We use 1-shot prompts in our main experiments and analyze the impact of the number of shots with 5- and 10-shot prompts separately in section 5.3).
@@ 4.2 Evaluation Metrics
We used the same set of metrics as in the CEVAL benchmark [22] to assess both counterfactual fidelity and text quality.
@@ 4.2.1 Counterfactual metrics
measures how effectively a method can change labels of instances with respect to a pretrained classifier $f$. FR is defined as the percentage of generated instances where the labels are flipped over the total number of instances $N$ [1]:
@@ 4.2.2 Text quality metrics
Text quality metrics are assessed on a 5-point scale by prompting GPT-3.5 with a temperature of 0.2 in line with the benchmark [22].
@@ 5.1 Incorporating Classifier Information Improves Performance on Counterfactual Metrics
Table 1 shows that our classifier-guided approaches improve key counterfactual metrics compared to the $vanilla$ baseline across both datasets and for all LLMs.
Table 1: Results for counterfactual metrics (FR: flip rate, Dis: token-level Levensthein Distance, PP: perplexity) and text quality metrics (GM: grammar, Coh: Cohesiveness, Flu: Fluency) on IMDB and SNLI. Italic denotes the best value within the same LLM, bold denotes the best value across all models, and the second-best value is underlined. vanilla indicates that no classifier information is used, and CGGV denotes the combination of CGG and CGV.
@@
For the combination of our two approaches (CGGV), we observe a consistent decrease of distance scores (Dis) and increase in flip rate (FR) in both datasets and across all LLMs. Among the language models in the combination of our two approaches (CGGV), Llama-3.1 achieves the best results, outperforming the flip rate of CREST and FLARE and closely matching MICE, the state-of-the-art counterfactual generation method in terms of flip rate. Notably, with more shots, Llama-3.1 even surpasses MICE in FR on the SNLI dataset (see Section 5.3). These results are particularly encouraging, as CREST, FLARE and MICE are dedicated counterfactual generation methods that require task-specific fine-tuning. In contrast, despite their simplicity, our approaches are equally (or even more) effective in generating high-fidelity counterfactuals. While text quality metrics (Grammar: GM, Fluency: Flu, Cohesiveness: Coh) show a slight decline – likely due to distribution shifts from forced word modifications – our approaches still outperform all other methods on these metrics except for FLARE on the IMDB dataset where performance is close. The high performance of CGV highlights that LLMs can generate high-quality counterfactuals with simple prompting, requiring only classifier-provided labels for guidance.
@@ 5.2 Not All LLMs Benefit Equally from Feature Importance Guidance Signals
We investigate whether LLMs effectively use the words identified as important by the XAI method in CGG by measuring the modification rate (MR). MR is the percentage of important words that were changed, i.e., words present in the original text but missing from the generated text. As shown in Fig. 3(a), CGG increases MR significantly from 0–40% to 70–75%. However, Fig. 3(b) shows that high MR does not guarantee flipping labels. While in LLama models, instances with flipped labels (true counterfactuals) are associated with higher MR than non-flipped instances, the opposite holds for GPT-4o-mini. Inspecting the flip rate for counterfactuals that rely heavily on important words (change at least 50% of them) shows that only Llama-2 benefits from this guidance signal, whereas flip rate decreases for the other two LLMs. This decrease could be the result of two potential reasons: (1) words that are important for the original prediction (label $y$) may not be relevant for counterfactuals, as noted in prior work [32], and (2) LLMs may fail to provide proper word substitutions.
Table 2 presents examples where LLMs succeed or fail to flip the label. In failure cases, despite changing all important words the label remained unchanged, because the words were deleted or replaced by synonyms. In successful cases, LLMs not only changed important words but also modified other words to improve coherence, resulting in better text quality.
@@
Figure 3: Impact of CGG on modification rate (MR) and flip rate (FR) in the SNLI dataset. Fig. 3(a) illustrates the MR for each LLM, with and without CGG. Fig. 3(b) compares the MR for flipped (Success) and non-flipped (Fail) instances, alongside the FR for cases where MR $>$ 0.5 and MR $\leq$ 0.5.

### v2

@@ Abstract
The need for interpretability in deep learning has driven interest in counterfactual explanations, which identify minimal changes to an instance that change a model’s prediction. Current counterfactual (CF) generation methods require task-specific fine-tuning and produce low-quality text. Large Language Models (LLMs), though effective for high-quality text generation, struggle with label-flipping counterfactuals (i.e., counterfactuals that change the prediction) without fine-tuning. We introduce two simple classifier-guided approaches to support counterfactual generation by LLMs, eliminating the need for fine-tuning while preserving the strengths of LLMs. Despite their simplicity, our methods outperform state-of-the-art counterfactual generation methods and are effective across different LLMs, highlighting the benefits of guiding counterfactual generation by LLMs with classifier information. We further show that data augmentation by our generated CFs can improve a classifier’s robustness. Our analysis reveals a critical issue in counterfactual generation by LLMs: LLMs rely on parametric knowledge rather than faithfully following the classifier. [^1] Source code is available at https://github.com/bach1292/llm-cls
@@ 1 Introduction
Counterfactual explanations identify minimal changes to the input that result in a changed prediction by a classifier. As such, they are a core tool for eXplainable AI (XAI), providing valuable insights into a classifier’s decision-making process [miller-2019-explanation, wu-etal-2021-polyjuice], thereby fostering interpretability. Counterfactual explanations in an XAI context differ from real-world counterfactuals: the latter describe hypothetical scenarios that contradict known facts [kulakova2013processing], while the former aim to uncover a classifier’s decision-making process. For example, given a negatively classified movie review, identifying minimal changes (adding, replacing or deleting words) that would turn the classifier’s prediction into positive, while maintaining meaningful content (cf. Fig.˜1) can provide insight into the classifier’s reasoning and decision boundary. This insight can expose undesired model behavior, such as biases against minorities [mehrabi2021survey], reliance on spurious correlations [geirhos2020shortcut], or vulnerability to adversarial manipulations [goodfellow2014explaining].
Counterfactual generation methods that access the classifier produce counterfactuals (CFs) with a high label-flip rate (i.e., they succeed in actually changing the prediction) and minimal changes [ross-etal-2021-explaining, treviso-etal-2023-crest, robeer-etal-2021-generating-realistic]. However, they require task-specific fine-tuning and may generate low-quality text [nguyen-etal-2024-ceval-benchmark]. In contrast, Large Language Models (LLMs) are used for end-to-end generation without expensive fine-tuning and produce high-quality text [radford2019language, kojima2022large]. However, they struggle to flip the labels with minimal changes [nguyen-etal-2024-ceval-benchmark, bach2024llms], i.e., do not generate high-fidelity [^2] The terms fidelity and faithful are sometimes used synonymously to describe how well an explanation aligns with the true reasoning of a classifier. In this paper, we use fidelity to refer to “true counterfactuals”, i.e., counterfactuals that actually change a classifier’s prediction and faithful to describe whether a counterfactual truly explains the classifier’s actual reasoning. CFs, because they do not have information about the classifier and its decision boundary.
@@
In this paper, we aim to bridge the gap between fully supervised, costly CF generation methods and unsupervised, end-to-end text generation using LLMs. We present two simple yet effective approaches for incorporating classifier information to guide LLMs in CF generation (cf. Fig.˜1), enabling high-quality text generation without the need for fine-tuning and offering broad applicability. In an ante-hoc approach, we first identify important words using an XAI method (saliency maps [simonyan2013deep] and SHAP [lundberg2017shap]). We then use the most important words and examples of factual-counterfactual pairs to prompt the LLM to generate counterfactual texts. Our second approach prompts the LLM to generate a set of candidate counterfactuals, evaluates their fidelity to the classifier, and selects the best candidate post-hoc based on minimality of changes and fidelity.
@@ 2 Related work
Counterfactual text generation methods fall into three categories [nguyen-etal-2024-ceval-benchmark]: masking and filling approaches, sampling based on conditional distributions, and LLM-based generation.
@@
These methods involve (1) masking key words and (2) filling the masked words using a pretrained language model. MICE [ross-etal-2021-explaining] and AutoCAD [wen-etal-2022-autocad] use classifier gradients to identify important words, while DoCoGen [calderon-etal-2022-docogen] masks n-grams with scores exceeding a threshold. CREST [treviso-etal-2023-crest] trains SPECTRA [guerreiro-martins-2021-spectra] to detect masking candidates. In step (2), these methods fine-tune T5 [raffel2020exploring] to fill the blanks. Polyjuice [wu-etal-2021-polyjuice] allows user-defined masking and fine-tunes a RoBERTa-based model using control codes to fill the blanks.
@@
CF-GAN [robeer-etal-2021-generating-realistic] and CLOSS [fern-pope-2021-text] model a conditional distribution to generate counterfactuals based on the target label. CF-GAN adapts StarGAN [choi2018stargan] to produce realistic counterfactuals. CLOSS generates text counterfactuals by optimizing token embeddings, ranking impactful changes with Shapley values, and using beam search to select minimal, grammatically plausible edits that flip a model’s prediction.
@@
CORE [dixit-etal-2022-core] trains a counterfactual retriever using contrastive learning, where positive pairs consist of human-authored counterfactuals and negative pairs are paraphrases. The retriever selects relevant counterfactual excerpts from an unlabeled corpus, which are then incorporated into prompts to guide GPT-3 [brown2020language] in generating diverse counterfactual edits. CORE relies on counterfactual data and does not use classifier information. In contrast, DISCO [chen-etal-2023-disco] generates a large set of perturbations by masking spans identified by a neural syntactic parser and filling them using GPT-3. A filtering model then selects the most promising counterfactuals. However, DISCO is computationally expensive and therefore primarily suited for short-text tasks such as natural language inference (NLI). Additionally, DISCO focuses on generating CFs for data augmentation rather than explanations. FLARE [bhattacharjee2024towards] generates CFs by prompting LLMs in three steps: extracting latent features, identifying relevant words linked to those features, and minimally modifying these words to produce CFs.
In summary, MF and CD methods require task-specific fine-tuning, while LLM-based generation lacks classifier integration and produces poor fidelity counterfactuals [nguyen-etal-2024-ceval-benchmark]. Our work not only addresses the limitations of both, fine-tuned and LLM-based methods, but also combines their strengths.
@@ 3.1 Classifier-Guided Generation (CGG)
In CGG, the generation process is guided by information from the classifier. Specifically, we use an XAI method $X(f(x))$ to extract words from the input $x$ that are important to the classifier’s decision $f(x)$. These important words are then incorporated into the predefined prompt $p$ to form $p(x,X(f(x)))$ (see Fig.˜2(a)). A complete example of the prompt is provided in Appendix˜0.A. This prompt is input to the LLM $M$, which generates a counterfactual instance $x^{\prime}$, such that $M(p(x,X(f(x))))=x^{\prime}$.
@@ 3.2 Classifier-Guided Validation (CGV)
CGV uses the classifier after the generation process. Specifically, we use the LLM $M$ with prompt $p$ to generate $n$ candidate samples $\{\hat{x}_{i}\}_{i=1}^{n}$. The classifier $f$ is then used to validate the labels, assigning a label $\hat{y}_{i}$ to each sample, and we select the sample with the flipped target label $y^{\prime}$ that also minimizes the distance $d(\hat{x}_{i},x)$ to the original input $x$ (see Fig.˜2(b)). Formally, we have
@@ 4 Experimental Setup
We evaluate counterfactuals on the CEVAL benchmark [nguyen-etal-2024-ceval-benchmark], which includes two datasets: IMDB [maas-etal-2011-learning], a sentiment analysis dataset with binary labels (positive and negative), and SNLI [bowman-etal-2015-large], a natural language inference dataset with three labels (entailment, contradiction, and neutral). The benchmark assesses performance by counterfactual and text quality metrics. A benchmark run required approximately two hours per dataset on an A100 80GB GPU.
@@ 4.1 Models
Following the benchmark [nguyen-etal-2024-ceval-benchmark], we use BERT-based classifiers $f$ that achieve 89% accuracy on IMDB and 90% on SNLI. We compare against MICE [ross-etal-2021-explaining], the top method for counterfactual metrics according to the benchmark, CREST [treviso-etal-2023-crest], a masking and filling method that uses SPECTRA [guerreiro-martins-2021-spectra] for masking and T5 [raffel2020exploring] for filling in the blanks, and FLARE [bhattacharjee2024towards], an LLM-based method that generates counterfactuals through three steps by prompting LLMs. The baseline is vanilla 1-shot prompting without classifier guidance.
For the counterfactual generation LLM $M$, we compare Llama-3.1-8B-Instruct, Llama-2-7b-chat, and GPT-4o-mini [^3] Version: GPT-4o-mini-2024-07-18.. Following [nguyen-etal-2024-ceval-benchmark], we use a temperature of 1.0 for all LLMs in the counterfactual generation task. For CGG, we adopt saliency maps [simonyan2013deep] as the XAI method $X$, selecting the top 25% most important words with the highest gradients relative to the classifier’s prediction, consistent with the benchmark (additional results with SHAP [lundberg2017shap] are similar, cf. Appendix˜0.B). For CGV, we use Levenshtein Distance as distance measure $d$ to select $x^{\prime}$ and set the number of samples $n=5$ for IMDB and $n=10$ for SNLI, as larger $n$ values showed minimal differences in preliminary experiments. We use 1-shot prompts in our main experiments and analyze the impact of the number of shots with 5- and 10-shot prompts separately in section 5.3).
@@ 4.2 Evaluation Metrics
We used the same set of metrics as in the CEVAL benchmark [nguyen-etal-2024-ceval-benchmark] to assess both counterfactual fidelity and text quality.
@@ 4.2.1 Counterfactual metrics
measures how effectively a method can change labels of instances with respect to a pretrained classifier $f$. FR is defined as the percentage of generated instances where the labels are flipped over the total number of instances $N$ [bhattacharjee2024towards]:
@@ 4.2.2 Text quality metrics
Text quality metrics are assessed on a 5-point scale by prompting GPT-3.5 with a temperature of 0.2 in line with the benchmark [nguyen-etal-2024-ceval-benchmark].
@@ 5.1 Incorporating Classifier Information Improves Performance on Counterfactual Metrics
Table 1 shows that our classifier-guided approaches improve key counterfactual metrics compared to the vanillavanilla baseline across both datasets and for all LLMs.
Table 1: Results for counterfactual metrics (FR: flip rate, Dis: token-level Levensthein Distance, $P$: perplexity) and text quality metrics (GM: grammar, Coh: Cohesiveness, Flu: Fluency) on IMDB and SNLI. Italic denotes the best value within the same LLM, bold denotes the best value across all models, and the second-best value is underlined. vanilla indicates that no classifier information is used, and CGGV denotes the combination of CGG and CGV.
@@
For the combination of our two approaches (CGGV), we observe a consistent decrease of distance scores (Dis) and increase in flip rate (FR) in both datasets and across all LLMs. Among the language models in the combination of our two approaches (CGGV), Llama-3.1 achieves the best results, outperforming the flip rate of CREST and FLARE and closely matching MICE, the state-of-the-art counterfactual generation method in terms of flip rate. Notably, with more shots, Llama-3.1 even surpasses MICE in FR on the SNLI dataset (see Section˜5.3). These results are particularly encouraging, as CREST, FLARE and MICE are dedicated counterfactual generation methods that require task-specific fine-tuning. In contrast, despite their simplicity, our approaches are equally (or even more) effective in generating high-fidelity counterfactuals. While text quality metrics (Grammar: GM, Fluency: Flu, Cohesiveness: Coh) show a slight decline – likely due to distribution shifts from forced word modifications – our approaches still outperform all other methods on these metrics except for FLARE on the IMDB dataset where performance is close. The high performance of CGV highlights that LLMs can generate high-quality counterfactuals with simple prompting, requiring only classifier-provided labels for guidance.
@@ 5.2 Not All LLMs Benefit Equally from Feature Importance Guidance Signals
We investigate whether LLMs effectively use the words identified as important by the XAI method in CGG by measuring the modification rate (MR). MR is the percentage of important words that were changed, i.e., words present in the original text but missing from the generated text. As shown in Fig.˜3(a), CGG increases MR significantly from 0–40% to 70–75%. However, Fig.˜3(b) shows that high MR does not guarantee flipping labels. While in LLama models, instances with flipped labels (true counterfactuals) are associated with higher MR than non-flipped instances, the opposite holds for GPT-4o-mini. Inspecting the flip rate for counterfactuals that rely heavily on important words (change at least 50% of them) shows that only Llama-2 benefits from this guidance signal, whereas flip rate decreases for the other two LLMs. This decrease could be the result of two potential reasons: (1) words that are important for the original prediction (label $y$) may not be relevant for counterfactuals, as noted in prior work [wu2021polyjuice], and (2) LLMs may fail to provide proper word substitutions.
Table˜2 presents examples where LLMs succeed or fail to flip the label. In failure cases, despite changing all important words the label remained unchanged, because the words were deleted or replaced by synonyms. In successful cases, LLMs not only changed important words but also modified other words to improve coherence, resulting in better text quality.
@@
Figure 3: Impact of CGG on modification rate (MR) and flip rate (FR) in the SNLI dataset. Fig.˜3(a) illustrates the MR for each LLM, with and without CGG. Fig.˜3(b) compares the MR for flipped (Success) and non-flipped (Fail) instances, alongside the FR for cases where MR >> 0.5 and MR $\leq$ 0.5.